A cylinder has a radius of 5 in. If the volume of the cylinder is 250x3.14 in.to the 3rd power, what is the height of the cylinder? : A cylinder has a radius of 5 in. If the volume of the cylinder is 250x3.14 in.to the 3rd power, what is the height of the cylinder?
Answer by checkley71(8405)

vol=pi*r^2*h
250pi=pi*5^2*h
250=25*h
h=250/25
h=10 inches.
proof
250*pi=pi*25*10
250pi=250pi